# Real Time Multi Organ Classification on Computed Tomography Images

### v1

@@ Abstract
Organ segmentation is a fundamental task in medical imaging, and it is useful for many clinical automation pipelines. Typically, the process involves segmenting the entire volume, which can be unnecessary when the points of interest are limited. In those cases, a classifier could be used instead of segmentation. However, there is an inherent trade-off between the context size and the speed of classifiers. To address this issue, we propose a new method that employs a data selection strategy with sparse sampling across a wide field of view without image resampling. This sparse sampling strategy makes it possible to classify voxels into multiple organs in real time without using accelerators. Although our method is an independent classifier, it can generate full segmentation by querying grid locations at any resolution. We have compared our method with existing segmentation techniques, demonstrating its potential for superior runtime in practical applications in medical imaging.
@@ 1 Introduction
Medical image segmentation has been a long-standing research topic within the healthcare and medical imaging communities, aimed at enhancing radiology workflows through automation. Compared to landmarking and bounding box detection methods, segmentation provides more granular information. This is beneficial for storing the anatomical locations of findings in structured databases, quantifying abnormalities, planning radiotherapy, calculating doses, comparing longitudinal studies, visualizing imaging data, or filtering the scanning region for CAD algorithms. Therefore, it is desirable to have fast and accurate segmentation algorithms in our medical vision pipelines.
@@
The field of medical image segmentation has significantly evolved, transitioning from traditional methodologies to advanced deep learning algorithms and thereby achieving improvements in both the quality and speed of segmentation processes. A notable development in this domain has been the introduction of U-net and its variants, which leverage feature pyramid structures to combine global and local information, enhancing robustness and achieving higher resolution masks [11]. This approach has not only proved beneficial for medical image segmentation but has also found applicability in a broad range of segmentation tasks across general computer vision.
@@
Recent advancements have introduced transformer architectures to overcome the limitations of convolutional networks’ receptive fields, thus enabling superior segmentation outcomes in medical images[9, 8]. These methods harness token-to-token interactions across the entire field of view within small patches to capture long-range information. However, the architecture change introduces quadratic complexity in relation to the field of view, posing challenges for computational efficiency.
@@
Independent classifiers are not commonly used for medical image segmentation because they are considered inefficient due to the repetitive computations required by the sliding window technique. Moreover, the absence of dependency between classification estimations in neighboring locations creates more irregular masks. However, they offer the advantage of quickly obtaining the organ label without needing to scan the entire input image. In this study, we demonstrate a method to reduce the computational burden of classifier-based segmentation using a data selection strategy. This strategy involves sparse sampling of intensities with a wide field of view in a hierarchical grid, similar to the approach of Yerebakan et al[18]. This selection strategy allows a machine learning classifier to label only query locations with a reduced amount of data while maintaining segmentation-level accuracy. Furthermore, applying the classifier to a regular grid on an image at any resolution provides full-volume segmentation. We also describe a refinement operation that improves the granularity of masks by increasing resolution at edges. With this architecture, we can avoid significant computation and storage costs, making many applications practical, such as storing all user annotated findings with corresponding organ labels or organ-specific processing of input images without the need for full segmentation. In our experiments, we used the public BTCV dataset for evaluation. Classifying a single point takes about 0.92 ms, whereas fine-level segmentation takes around 10 s without requiring additional GPU hardware.
@@ 1.1 Related Work
Identifying organs for points within medical images can be approached through various anatomical intelligence methods. For example, landmarking techniques offer quick estimates of anatomical regions but have limited precision for specific organ labeling at arbitrary positions [5]. Similarly, object detection methods provide bounding box definitions but encounter limitations in precise organ boundaries. Atlas-based segmentation can accurately label organs, involving image registration to a predefined atlas. Thus, this approach is not preferred due to its high computational demands and the necessity of maintaining atlas images alongside their segmentations, making it less practical for real-time applications. Even strategies focusing on semantic feature matching to match points to an atlas [18, 2, 16] fail to achieve processing times under 100ms, indicating a gap in achieving real-time performance.
Organ classification, while less extensively explored than segmentation techniques, finds significant application in the MedMNIST dataset [17]. Although MedMNIST serves as an effective educational benchmark, its utility in real-world scenarios is limited due to varying views of organ fields, necessitating object detection to standardize image sizes for direct application. For that, Object detection, followed by classification, forms a pipeline approach but requires additional complexity. Xu et al. [15] demonstrated that organ bounding box detection of organs could be achieved in just 0.25 seconds using acceleration hardware. However, bounding boxes provide only a rough estimation and cannot pinpoint specific organ locations based on user interactions, such as a mouse click position. In contrast, organ segmentation offers a more detailed view than simple landmark detectors or bounding boxes, providing the granularity necessary for precise organ identification. Consequently, our research primarily focuses on exploring the alternatives of segmentation methods and specifically addressing the need for precomputation, which is costly.
@@
U-Net, introduced by Ronneberger et al. [11, 6], has become a baseline for segmentation algorithms for a considerable time. This architecture is initially designed for medical image segmentation using CNN architectures. U-Net’s strength lies in its utilization of multiple image resolutions to extract feature vectors, allowing coarse layers to focus on global representations while finer levels handle local details. Several studies have modified various aspects of U-Net to enhance segmentation performance[19]. In our method, we also employ a similar multi-resolution strategy with a large field of view. However, instead of resampling the image volume, we have directly defined a sampler function to select intensities from the raw data with offset locations. This sampling mechanism allows the creation of independent descriptors at voxel locations, allowing rapid point-wise classifications.
@@
Recently, transformers have gained popularity in the medical imaging domain. UnetTr [9], in particular, has demonstrated superior performance compared to U-Net on the BTCV dataset. Subsequent improvements have addressed both computational efficiency and accuracy. For instance, SwinUnetTR [3] enhanced the UnetTr framework and shifted window strides in segmentation tasks. Despite these advancements, the computational time for these algorithms remains a challenge, especially for real-time applications, even when utilizing GPU hardware.
@@
Foundational model approaches have gained popularity in medical vision, similar to Large Language Models in NLP. An example is Vox2Vec [7], which introduces an alternative method for learning directly from the image data without supervision. There are studies that also addressed pretraining self-supervised models for better segmentation accuracy [4]. These representation learning methods can create vectors for individual voxels that could be used for classification. Still, processing a large subvolume is necessary, even for a single point.
It is noteworthy that the dataset space for medical image segmentation tasks is also expanding. The Total Segmentation Dataset has published 1,204 images for 117 organs [14]. The model has been made available within the Monai [^1] https://monai.io/ library and can be used directly as a command-line program. The algorithm’s runtime is approximately 30 seconds on GPU and several minutes on CPU hardware. Additionally, there are various competitions in medical image segmentation. Among them, the Flare challenge aims for fast computation times, with the fastest methods clocking around 10 seconds using GPU hardware  [^2] https://flare22.grand-challenge.org/testing-results/[10] in the 2022 leaderboard.
@@ 2 Our Method
(a) Sampling Offsets For Descriptor Definition
(b) Decoding Image From A Descriptor
Figure 1: Descriptor Definition
@@ 2.1 Sparse Sampling Image Intensities
The resampling step is very common in medical image segmentation pipelines. It remains time-consuming for interactive use and requires additional memory. Instead of resampling the image, our method defines a sampler function with fixed millimeter offsets. Then, descriptors are just image intensities on those offset locations. This sampler works only on selected query points, significantly reducing the required computation for any point of interest in the image.
@@
The organ label in a point of interest depends on the intensity values near the point itself. On the other hand, context plays a role in disambiguating similar-looking soft tissue types. Therefore, instead of sampling with a single resolution and a limited subvolume, we have used different regular grids at different resolutions for sampling to capture a large area in a hierarchical manner. Thanks to the consistency of the human body, similar anatomical locations generate similar descriptors. The fixed offsets behave like distance sensors that encode location.
@@
The sampler offsets are in millimeters, and they are adjusted to voxel offsets once the image is loaded, using division by voxel spacing and rounding. Also, if the offset locations are outside the image volume, they are assigned a value of 0 for the corresponding dimension, ensuring all descriptors are in the same vector space. During runtime, descriptor computation involves only memory lookup operations. Memory positions come from adding offsets to the current query point. In our experiments, we have utilized 2D+3D grids for sampling. The 2D part consists of 3 orthogonal planes defined at a 4mm resolution by a 27x27 grid. The 3D part comprises six three-dimensional grids at resolutions of 2, 3, 5, 12, 28, and 64 mm, respectively, with 9x9x9 grid sizes considering some slice thickness in CT images. Spacing increased with non integer multipliers to avoid sampling the same locations in different resolution grids. Selection of size 9 grid allows both 2d and 3d to be present in the same picture simplifying debugging. We named this approach 3.5D sampling since it contains more than 3D information. Exemplary sampling grids in a single slice given a point are shown in Figure 1(a), where bright dots are the sampling locations. The hierarchical nature of the sampling allows for a very large field of view with granular detail in the center. These descriptors can be decoded back into an 81x81 2D image as seen in Figure 1(b) by placing every 27x27 block in 9 different places. In this figure, the upper left corner is reserved for the 2D portion of the descriptor with three orthogonal planes, and the other parts are 3D slices at different resolutions with 9x9x9 grids. The total dimension of the descriptor becomes 6561. As seen in the figure, the query location belongs to the heart. While other descriptor definitions are possible, for the sake of simple decoding for visualization, we used this 9x9x9x9 descriptor definition described here. The organ label is defined as the segmentation label at the query point.
@@ 2.3 Segmentation
The classifier provides a functionality to get an organ label for any query location. Thus, it is possible to compute a segmentation mask by independently querying all voxels. However, querying every location in the image is computationally expensive. Instead, we can use a sparse grid with any resolution to obtain faster results.
(a) Coarse Segmentation
(b) Edge Refinement
(c) Fine Details
Figure 3: Classifier based segmentation allows to create segmentation masks in different resolutions without need of any change
The high-level 8mm sparse grid would create a coarse segmentation mask, as shown in an example in Figure 3(a) where each intensity represents a different organ label. Once this mask is generated, the whole volume could be filled with nearby labels. However, masks are not accurate near edges due to low resolution.
Since errors are happening on the edges, we could further refine the resolution in those regions in a hierarchical way. To achieve this, we have checked all points in the segmentation mask with higher resolution and added points into the task queue for the query if the neighbors have different labels, as shown in Figure 3(b). For each level, we have some smoothing on the edges by thresholding a majority of 20 within 27 neighbors to prevent excessive computation. Thus, the cost of classification in high resolution at nearly homogeneous points would be eliminated. We reduced spacing to 2mm with this approach. An example generated mask is shown in Figure 3(c). Thanks to points being completely independent of each other, there is a massive parallelization potential.
@@ 3 Experiments
(a) Image Slice
(b) Ground Truth Mask
(c) Coarse Segmentation
(d) Fine Mask
@@
We have experimented with our method on the public BTCV dataset. It contains a total of 30 images with 13 abdominal organ labels. We trained with 24 cases and used the validation part of the official split, which consists of 6 images for evaluation similar to compare studies. The data is accessible online  [^3] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789.
@@
Once the model is trained, it is converted to C++ counterpart to reduce latency. We have used OpenMP parallelization for different query points for segmentation part of the experiment. Time has measured on 6 core - 12 thread CPU.
@@
In the segmentation part, we first obtained coarse segmentation for full volume. An example 2d slice is shown in Figure 4(b). The average coarse segmentation speed is 5.24±0.9s per volume. Segmentation is also refined in the edges using the same classifier. The speed, in this case, is an average of 9.51±2.72s on the CPU. We measured dice scores and compared them against recently published studies on the same set. In terms of runtime speed, Shaker et. al. [12] reports GPU time of 62.4 seconds while being the most efficient algorithm therein. The results are shown in Table 1.
Table 1: Dice Score Averages on Multi Organ Segmentation on BTCV validation set
Methods Spleen RKid LKid Gall Eso Liv Sto Aor IVC Veins Pan RAG LAG AVG
From Scratch
ViT3D-B [4] 0.8902 0.8926 0.8769 0.4763 0.4891 0.9447 0.7475 0.8207 0.773 0.6175 0.6442 0.5663 0.4699 0.7084
nnFormer [12] 0.9458 0.8862 0.9368 0.6529 0.7622 0.9617 0.8359 0.8909 0.8080 0.7597 0.7787 0.7020 0.6605 0.8162
UnetTR++ [12] 0.9494 0.9190 0.9362 0.7075 0.7718 0.9595 0.8515 0.8928 0.8314 0.7691 0.7742 0.7256 0.6817 0.8328
Pretrained
ViT3D-L[4] 0.9556 0.9582 0.9414 0.5206 0.5352 0.9898 0.8025 0.8811 0.8298 0.6649 0.6916 0.6088 0.5045 0.7603
Ours 0.8398 0.8364 0.7990 0.4780 0.6489 0.9264 0.6860 0.7770 0.7106 0.5007 0.5583 0.4054 0.4089 0.6596
@@ 4 Discussion
The independence of the classifier is a limitation in segmentation operation. It is possible to create an additional filter on extracted class probabilities further to reduce the variance within local estimations. Another alternative is utilizing independent classifier features as token embedding for further transformer layers instead of processing full 3D raw data. However, in this case, real-time response needs to be sacrificed. The power of an independent classifier is that it works on any resolution and any interest region without the need for full-volume processing.

### v2

# cvprbluergb0.21,0.49,0.74

@@ Abstract
Organ segmentation is a fundamental task in medical imaging since it is useful for many clinical automation pipelines. However, some tasks do not require full segmentation. Instead, a classifier can identify the selected organ without segmenting the entire volume. In this study, we demonstrate a classifier based method to obtain organ labels in real time by using a large context size with a sparse data sampling strategy. Although our method operates as an independent classifier at query locations, it can generate full segmentations by querying grid locations at any resolution, offering faster performance than segmentation algorithms. We compared our method with existing segmentation techniques, demonstrating its superior runtime potential for practical applications in medical imaging.
@@ 1 Introduction
Medical image segmentation has been a long-standing research topic within the healthcare and medical imaging communities, aimed at enhancing radiology workflows through automation. Compared to landmarking and bounding box detection methods, segmentation provides more granular information. This is beneficial for storing the anatomical locations of findings in structured databases, quantifying abnormalities, radiotherapy planning, calculating doses, comparing longitudinal studies, visualizing imaging data, or filtering the operating region for CAD algorithms. Therefore, it is desirable to have fast and accurate segmentation algorithms in our medical vision pipelines.
@@
The field of medical image segmentation has significantly evolved, transitioning from traditional methodologies to advanced deep learning algorithms and thereby achieving improvements in both the quality and speed of segmentation processes. A notable development in this domain has been the introduction of U-net and its variants, which leverage feature pyramid structures to combine global and local information, enhancing robustness and achieving higher resolution masks [ronneberger2015u]. This approach has not only proved beneficial for medical image segmentation but has also found applicability in a broad range of segmentation tasks across general computer vision.
@@
Recent advancements have introduced transformer architectures to overcome the limitations of convolutional networks’ receptive fields, thus enabling superior segmentation outcomes in medical images[hatamizadeh2022unetr, hatamizadeh2021swin]. These methods harness token-to-token interactions across the entire field of view within small patches to capture long-range information. However, the architecture change introduces quadratic complexity in relation to the field of view, posing challenges for computational efficiency.
@@
Independent classifiers are not commonly used for medical image segmentation due to the inefficiency of repetitive sliding window computations and the irregular masks they produce. However, they can quickly obtain organ labels without scanning the entire image. We demonstrate a method to reduce the computational burden of classifier-based segmentation using a data engineering strategy that involves sparse sampling of intensities with a wide field of view in a hierarchical grid, similar to Yerebakan et al. [pointmatching]. This approach allows a classifier to label query locations with less data while maintaining segmentation level accuracy. Applying the classifier to a regular grid at any resolution provides full-volume segmentation, and a refinement operation improves mask granularity at edges. This architecture avoids significant computation and storage costs, making practical applications like storing annotated findings with organ labels feasible. Our experiments on the BTCV dataset show that classifying a single point takes about 0.92 ms, and fine-level segmentation takes around 10 s without additional GPU hardware.
@@ 1.1 Related Work
Identifying organs in medical images can be approached through various methods. Landmarking techniques provide quick estimates but lack precision for specific organ labeling at arbitrary positions [ghesu2017robust]. Object detection methods define bounding boxes but struggle with precise boundaries. Atlas-based segmentation will be robust thanks to registration but computationally demanding and impractical for real-time applications. Even semantic feature matching strategies fail to achieve real-time processing under 100ms [pointmatching, bai2023sam++, yan2022sam].
@@
U-Net, introduced by Ronneberger et al. [ronneberger2015u, gibson2018automatic], has long been a baseline for segmentation algorithms, utilizing multiple image resolutions to balance global and local details. Various studies have modified U-Net to enhance performance [zhou2018unet++]. Our method employs a similar multi-resolution strategy but uses a sampler function to select raw data intensities. This sampling mechanism allows the creation of independent descriptors at voxel locations, allowing rapid point-wise classifications without resampling the image volume.
@@
Recently, transformers have gained popularity in medical imaging. UnetTr [hatamizadeh2022unetr] outperformed U-Net on the BTCV dataset, with further improvements in efficiency and accuracy seen in SwinUnetTR [cao2022swin]. Despite these advancements, computational time remains a challenge for real-time applications, even with GPU hardware.
@@
Foundational models, similar to Large Language Models in NLP, are gaining traction in medical vision. Vox2Vec [goncharov2023vox2vec] offers unsupervised learning from image data, while other studies focus on pretraining self-supervised models for better segmentation accuracy [chen2023masked]. These methods create voxel-level vectors for classification, but still require processing large subvolumes even for a single point.
The datasets and pretrained models for medical image segmentation are expanding. One notable dataset for classification is MedMNIST [yang2023medmnist], which serves as an educational benchmark. However, its real-world application is limited due to the varying size of regions of interest. Another notable dataset, the Total Segmentator Dataset, includes 1,204 images of 117 organs [wasserthal2023totalsegmentator], and the model is available in the Monai library. The algorithm runs in about 30 seconds on a GPU and several minutes on a CPU. Competitions like the Flare challenge aim for fast segmentation, with top methods reported to achieve around 10 seconds runtime on a GPU [myronenko2023automated]. This demonstrates the gap in utilizing segmentation methods for user-interfacing medical imaging applications.
@@ 2.1 Sparse Sampling Image Intensities
{subfigure}
[…]
Figure 1: Sampling Offsets For Descriptor
Figure 2: Decoding Image From A Descriptor
Figure 3: Descriptor Definition and Decoding
@@
The resampling step is very common in medical image segmentation pipelines. Despite being basic, it is still time-consuming for interactive use and requires extra memory. Instead of resampling the image, our method defines a sampler function with fixed millimeter offsets. Then, descriptors are simply the image intensities at those offset locations. This sampler works only on selected query points, significantly reducing the required computation for any point of interest in the image.
@@
The organ label at a point of interest depends on the intensity values near the point itself. Additionally, context is crucial for disambiguating similar-looking soft tissue types. Therefore, instead of sampling with a single resolution and a limited subvolume, we use different regular grids at various resolutions to capture a large area hierarchically. Thanks to the consistency of the human body, similar anatomical locations generate similar descriptors. The fixed offsets function like distance sensors, encoding location.
The sampler offsets are in millimeters and are adjusted to voxel offsets once the image is loaded, by dividing by voxel spacing and rounding. If the offset locations are outside the image volume, they are assigned a value of 0 for the corresponding dimension, ensuring all descriptors are in the same vector space. During runtime, descriptor computation involves only memory lookup operations. Memory positions are determined by adding offsets to the current voxel.
In our experiments, we utilized 2D+3D grids for sampling. The 2D part consists of 3 orthogonal planes defined at a 4mm resolution by a 27x27 grid. The 3D part comprises six three-dimensional grids at resolutions of 2, 3, 5, 12, 28, and 64 mm, respectively, with 9x9x9 grid sizes, considering some slice thicknesses in CT images. The spacing increases with non-integer multipliers to avoid sampling the same locations in different resolution grids. The selection of a size 9 grid allows both 2D and 3D to be present in the same picture, simplifying debugging. We named this approach 3.5D sampling since it contains more than 3D information.
@@
Exemplary sampling grids in a single slice given a point are shown in Figure 3, where bright dots are the sampling locations. The hierarchical nature of the sampling allows for a very large field of view with granular detail in the center. These descriptors can be decoded back into an 81x81 2D image as seen in Figure 3, by placing every 27x27 block in 9 different places. In this figure, the upper left corner is reserved for the 2D portion of the descriptor with three orthogonal planes, and the other parts are 3D slices at different resolutions with 9x9x9 grids. The total dimension of the descriptor becomes 6561. As seen in the figure, the query location belongs to the heart.
While other descriptor definitions are possible, for the sake of simple decoding for visualization, we used this 9x9x9x9 descriptor definition described here. The organ label is defined as the segmentation label at the query point.
@@ 2.3 Segmentation
The classifier provides functionality to obtain the organ label for any query location. Thus, it is possible to compute a segmentation mask by independently querying all voxels. However, querying every location in the image is computationally expensive. Instead, we can use a hierarchical grid with multiple resolutions to obtain faster results.
{subfigure}
[…]
Figure 5: Coarse Segmentation
Figure 6: Edge Refinement
Figure 7: Fine Details
Figure 8: Classifier based segmentation enables segmentation in any resolution thus allowing multiple steps of refinement for fine segmentation
We used a high-level 8mm sparse grid to create a coarse segmentation mask by filling the 8mm subvolumes with predicted labels as shown in an example in Figure 8 where each intensity represents a different organ label. However, masks are not accurate near edges due to low resolution. Since errors are happening on the edges, we could further refine the resolution in those regions hierarchically. To achieve this, we have checked all points in the segmentation mask with higher resolution grid and added points into the task queue to query if the neighbors have different labels, as shown in Figure 8. For each level, we have some smoothing on the edges by thresholding a majority of 20 within 27 neighbors to prevent excessive computation. Thus, the cost of classification in high resolution at nearly homogeneous points would be eliminated. We reduced the spacing to 2mm with this approach.
An example generated mask is shown in Figure 8. Thanks to points being completely independent of each other, there is a massive parallelization potential.
@@ 3 Experiments
{subfigure}
[…]
Figure 9: Image Slice
Figure 10: Ground Truth Mask
{subfigure}
[…]
Figure 11: Coarse Segmentation
Figure 12: Fine Mask
@@
We have experimented with our method on the public BTCV dataset, which contains a total of 30 images with 13 abdominal organ labels. We trained with 24 cases and used 6 cases for validation, following the official split for evaluation, similar to other studies. The data is accessible online  [^1] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789.
@@
Once the model is trained, it is converted to C++ counterpart to reduce latency. We have used OpenMP parallelization for different query points for segmentation part of the experiment. Time has measured on 6 core - 12 thread Xeon CPU.
@@
In the segmentation part, we first obtained coarse segmentation for full volume. An example 2D slice is shown in Figure 13. The average coarse segmentation speed is 5.24±0.9s per volume. Segmentation is also refined in the edges using the same classifier. The speed, in this case, is an average of 9.51±2.72s on the CPU. We measured dice scores and compared them against recently published studies on the same set. In terms of runtime speed, Shaker et. al. [shaker2022unetr++] reports GPU time of 62.4 seconds while being the most efficient algorithm therein. The results are shown in Table 1.
Table 1: Dice Score Averages on Multi Organ Segmentation on BTCV validation
[…]
! Methods Spleen RKid LKid Gall Eso Liv Sto Aor IVC Veins Pan RAG LAG AVG From Scratch ViT3D-B [chen2023masked] 0.8902 0.8926 0.8769 0.4763 0.4891 0.9447 0.7475 0.8207 0.773 0.6175 0.6442 0.5663 0.4699 0.7084 nnFormer [shaker2022unetr++] 0.9458 0.8862 0.9368 0.6529 0.7622 0.9617 0.8359 0.8909 0.8080 0.7597 0.7787 0.7020 0.6605 0.8162 UnetTR++ [shaker2022unetr++] 0.9494 0.9190 0.9362 0.7075 0.7718 0.9595 0.8515 0.8928 0.8314 0.7691 0.7742 0.7256 0.6817 0.8328 Pretrained ViT3D-L[chen2023masked] 0.9556 0.9582 0.9414 0.5206 0.5352 0.9898 0.8025 0.8811 0.8298 0.6649 0.6916 0.6088 0.5045 0.7603 Ours 0.8398 0.8364 0.7990 0.4780 0.6489 0.9264 0.6860 0.7770 0.7106 0.5007 0.5583 0.4054 0.4089 0.6596
@@ 4 Discussion
The independence of the classifier is a limitation in segmentation operation. It is possible to create an additional filter on extracted class probabilities further to reduce the variance within local estimations. Another alternative is utilizing independent classifier features as token embeddings for further transformer layers instead of processing full 3D raw data. However, in this case, real-time response needs to be sacrificed. The power of an independent classifier is that it works on any resolution and any region of interest without the need for full-volume processing.